OpenStudy (danieke):

help To rent a certain meeting room, a college charges a reservation fee of $48 and an additional fee of $7.90 per hour. The math club wants to spend less than $103.30 on renting the meeting room.What are the possible amounts of time for which they could rent the meeting room? Use t for the number of hours the meeting room is rented, and solve your inequality for t.

OpenStudy (janu16):

Im going yo use different nunvers so i am gonna use 43 instead of 48, 6.80 instead of 7.90, 103.30 instead of 77 Try solving it with your numbers C = 6.80t + 43 77 = 6.80t + 43 34 = 6.80t 34/6.8 = t = 5hr

OpenStudy (janu16):

I meant 77 instaed of 103.30

OpenStudy (janu16):

What do you get when you plug your nunbers?

OpenStudy (janu16):

@danieke